OpenStudy (anonymous):

(2/(x-2))-(8/(x^2-4))

OpenStudy (anonymous):

expand x2 -4 as (x+2)(x-2)

OpenStudy (anonymous):

if someone could show me the steps on how to cross multiply and simplify this i would appreciate it

OpenStudy (e.mccormick):

So, \[\frac{2}{x-2}-\frac{8}{x^2-4}\] Yah, start lik what Koel said. Then get a common denominator, do the subtraction, and finish. Because it is a negative, not an equals, it is not really a cross multiply.

OpenStudy (anonymous):

2/x-2 - 8/x2 -4 =2/x-2 -8/(x+2)(x-2) multiply the numerator and denominator of the first term with x+2(because that's what is missing) =[2(x+2) -8 ]/(x+2)(x-2) =(2x+4-8)/(x+2)(x-2) =2x-4/(x+2)(x-2) =2(x-2)/(x+2)(x-2) =2/x+2

OpenStudy (anonymous):

thank you so much! took me a while to see it but i finally understand it now. thank you